Based on the reaction below. How many grams of water would be form from 0.33 moles of C4H10 reacting with an unlimited amount of oxygen gas? (3pts) C4H10 + 6O2 --> 4CO2 + 5H2O

1
Expert's answer
2021-11-17T08:46:02-0500

Mole ratio C4H10:H2O

= 1:5

= 0.33×5= 1.65moles

Molar mass of H2O= 18.02

= 18.02×1.65=29.733g